# BiVLC: Extending Vision-Language Compositionality Evaluation with Text-to-Image Retrieval

### v1

@@ 5.2 TROHN-Img: Training on Hard Negative Images
GNM [Sahin et al., 2024] used automatically generated images as hard negative examples, but to generate them, GNM uses image editing techniques and is limited to the Replace category, i.e. given an original image, they only get hard negative images replacing one of the objects of the original image [Sahin et al., 2024]. That means that GNM is limited to the Replace category, ignoring Swap and Add categories.
To overcome those limitations, we design a new dataset named TROHN-Img (TRaining On Hard Negative Images). This dataset is based on the negative captions previously generated for TROHN-Text. Due to the resources and time needed for image generation, TROHN-Text is too big (>3M hard negative captions), so we decided to filter the instances to keep the best ones and obtain a training set equivalent to COCO in size [Lin et al., 2014]. This way, we can also measure better the contribution of the hard negative images, since the size of the training set is similar to the one used for our CLIPCOCO model, the best model for BiVLC so far. Thus, we apply the following steps to generate TROHN-Img:
@@ 5.3 Results and analysis
Table 3: Results on SugarCrepe and BiVLC, including our models (bottom rows). Bold for best, underline for second best. For BiVLC, we show the three main scores plus individual retrieval task scores: Ipos2T refers to image-to-text retrieval with positive image, Ineg2T for image-to-text retrieval with negative image, etc.
@@
Table 3 shows the results of our two proposed models, i.e. CLIPTROHN-Text and CLIPTROHN-Img compared to the other multimodal models, both in SugarCrepe and BiVLC  main metrics (leftmost four columns). As it can be observed, CLIPTROHN-Img outperforms all the other multimodal models by large margins for the main three metrics of BiVLC. Interestingly, the improvement comes mainly from text-to-image retrieval, which greatly benefits from training with hard negative images. However, CLIPTROHN-Img is still much better for image-to-text retrieval and its group score is well below human performance, which shows that further research is needed for achieving human comparable performance for bidirectional VLC.
On the other hand, CLIPTROHN-Text is the best model for SugarCrepe, outperforming NegCLIP by 8 points and closing the gap with humans to 5 points. However, this strong performance does not reflect on BiVLC, where CLIPTROHN-Text is even worse than CLIPCOCO, NegCLIP and GNM. This highlights our previous finding about the lack of correlation between SugarCrepe and BiVLC results (Section 4).
To put it into the context of the state-of-the-art, the strongest models for SugarCrepe are CapPa [Tschannen et al., 2024] with 92.88 and GPT-4V [Achiam et al., 2023] with 92.19111Results obtained from https://github.com/RAIVNLab/sugar-crepe/tree/main/gpt-4v-results, both private models. CLIPTROHN-Text outperforms both models, showing it is very strong for image-to-text retrieval. On the other hand, CLIPTROHN-Img is also strong, but it performs around 2.5 points worse than GPT-4V in SugarCrepe and 4 points worse than CLIPTROHN-Text. The lower performance of CLIPTROHN-Img on SugarCrepe can be attributed to TROHN-Img containing 10 times less hard negative captions than TROHN-Text, and could improve with more instances.
The complete results of all multimodal models in BiVLC according to caption generation categories are depicted in Table 6 (Appendix D), where the Swap category is the hardest one for models.
@@
We can also evaluate the detectors on the BiVLC retrieval tasks: (i) for image-to-text retrieval, we first use the visual encoder to predict the type of image (natural or synthetic); if the detection is correct, the output of the text detector is used to see which of the captions has the highest probability to be of the same type as the image and we select it as the matching caption; (ii) for text-to-image retrieval, we follow the same procedure, but using first the text detector over the given caption, and afterwards the visual detector for the two candidate images. The rightmost columns in Figure 4 show that BiVLC group score is very low for both detectors, underperforming clearly the multimodal models (cf. Table 3). This means that our models are actually learning much more than just distinguishing synthetic and natural images-captions. We already had another evidence of that in Table 3, where all models do equally well on Ipos2T and Ineg2T, i.e. they work equally well for both natural and synthetic images.
@@ 6 Conclusions
We have presented BiVLC, a dataset for bidirectional vision-language compositionality. Our dataset offers the possibility of evaluating multimodal models for image-to-text and text-to-image retrieval. We have shown that including text-to-image retrieval leads to novel findings, such as the unbalanced performance of multimodal models in the two retrieval directions, in contrast to humans. We also uncovered the low performance for BiVLC of the models which have been specifically trained for image-to-text retrieval. Furthermore, we have proposed a new way to train models with hard negative images, resulting on the best model for BiVLC so far. The gap with human performance shows that BiVLC can be used to measure further advances in modelling VLC.
@@ A.1 Limitations
BiVLC offers captions only in English. It would be interesting to extend the dataset to other languages, as some recent works in vision-language models are already doing [Pouget et al., 2024, Chen et al., 2023, Bugliarello et al., 2022]. Moreover, we only trained and evaluated contrastive models, due to their suitability for image-to-text retrieval tasks and their availability. In the future, other multimodal models could also be explored. Finally, as we rely on SugarCrepe hard negative captions, we also use the same categories. Adding more diversity by extending BiVLC to other categories could be beneficial.
@@ A.2 Societal impacts
Appendix B TROHN-Text dataset details
@@
Appendix C Detailed evaluation metrics
@@
Appendix D BiVLC results per category
BiVLC is formed by the same categories presented in SugarCrepe. This provides the possibility to analyze the performance of the models on different types of negatives. Table 6 presents the performance of the multimodal models for BiVLC, divided into the categories Replace, Swap and Add.
Table 6: Performance of multimodal models for BiVLC, divided into Replace, Swap and Add categories.
Model Replace Swap Add
I2T T2I Group I2T T2I Group I2T T2I Group
CLIP 82.09 60.36 57.27 46.52 16.16 13.65 70.32 44.63 39.58
CLIPCOCO 87.99 72.42 69.94 51.53 25.07 20.89 83.16 55.58 51.58
NegCLIP 86.37 70.80 68.03 47.08 24.23 18.66 81.26 51.37 48.00
GNM 86.14 68.70 65.75 50.70 17.83 15.88 83.16 58.74 55.37
CLIPTROHN-Text 83.61 70.03 65.84 54.32 25.35 20.61 72.21 55.37 48.42
CLIPTROHN-Img 91.62 79.18 76.61 62.40 32.03 27.86 94.74 69.47 68.00
@@ E.3 Software information
Fine-tuning hyper-parameters
@@
We base CLIP fine-tuning on OpenCLIP [Cherti et al., 2022]. Detailed hyper-parameters:
@@
We also trained CLIPDet and CLIPTROHN-Img/Det detectors for binary classification by keeping the encoders frozen and adding a sigmoid neuron over the CLS embedding for the image encoder and over the EOT embedding for the text encoder. Detailed hyper-parameters:
@@
  • Epochs: We trained the text detector over 10 epochs and the image detectors over 1 epoch. We used validation accuracy as the model selection criterion, i.e. we selected the model with highest accuracy in the corresponding validation set.
@@
We base our evaluation on OpenCLIP [Cherti et al., 2022]. We follow all the default hyperparameters used to evaluate models, making sure that when loading the checkpoints we are using QuickGELU as in the base pretrained model [^15] Evaluation bug when using GELU vs QuickGELU https://github.com/RAIVNLab/sugar-crepe/issues/7.
@@ Appendix F SugarCrepe templates
Appendix G BiVLC dataset information
Dataset documentation
  • image: COCO 2017 validation image.
  • caption: COCO 2017 validation text describing the COCO image.
  • negative_caption: Negative caption generated from the COCO text description by SugarCrepe.
  • negative_image: Negative image generated from the negative caption by BiVLC.
  • type: Category of the negative instances: Replace, Swap or Add.
Figure 4: Example of a BiVLC instance after loading the dataset.
Maintenance plan
Licensing
Author statement
Appendix H BiVLC instance examples
@@ Appendix I Crowdsourcing
I.1 Instructions for the different stages
@@
The instructions for this stage are divided into two: image-to-text retrieval (I2T) and text-to-image retrieval (T2I). Both are detailed below, and examples of question as seen by the annotators are depicted in Figures 8 and 9. The obtained human performance for each task can be seen in Table 2. {spverbatim} "Each survey consists of 100 questions/instances. We split the instances into 2 parts (50/50), in which the task changes.
@@
I.2 Hourly wage paid to participants and total amount spent
@@
I.3 Inter-tagger agreement

### v2

@@ 5.2 TROHN-Img: Training on Hard Negative Images
GNM [Sahin et al., 2024] used automatically generated images as hard negative examples, but to generate them, GNM uses image editing techniques and is limited to the Replace category. SyViC [Cascante-Bonilla et al., 2023] also explores training with synthetic images, but their images are generated with a graphic engine and require techniques such as style transfer to make them more natural.
To overcome those limitations, we design a new dataset named TROHN-Img (TRaining On Hard Negative Images). This dataset is based on the negative captions previously generated for TROHN-Text. Due to the resources and time needed for image generation, TROHN-Text is too big (>3M hard negative captions), so we decided to filter the instances to keep the best ones and obtain a training set equivalent to COCO in size [Lin et al., 2014]. This way, we can also measure better the contribution of the hard negative images, since the size of the training set is similar to the one used for our CLIPCOCO model, the best contrastive model for BiVLC so far. Thus, we apply the following steps to generate TROHN-Img:
@@ 5.3 Results and analysis
Table 3: Results for contrastive models of the same size on SugarCrepe and BiVLC, including our models (bottom rows). Bold for best, underline for second best. For BiVLC, we show the three main scores plus individual retrieval task scores: Ipos2T refers to image-to-text retrieval with positive image, Ineg2T for image-to-text retrieval with negative image, etc.
@@
Table 3 shows the results of our two proposed models, i.e. CLIPTROHN-Text and CLIPTROHN-Img compared to comparable contrastive models of the same size, both in SugarCrepe and BiVLC  main metrics (leftmost four columns). As it can be observed, CLIPTROHN-Img outperforms all the other contrastive models by large margins for the main three metrics of BiVLC. This shows the effectiveness of our training process. Interestingly, the improvement comes mainly from text-to-image retrieval, which greatly benefits from training with hard negative images. However, CLIPTROHN-Img is still much better for image-to-text retrieval and its group score is well below human performance, which shows that further research is needed for achieving human-comparable performance for bidirectional VLC. As mentioned in Section 4, we were able to run VQAScore and the open version of CapPa on BiVLC (see Table 2). CLIPTROHN-Img compares favorably to Open CapPa by a large margin, and it is comparable to VQAScore-XL, even though being orders of magnitude smaller (3B vs 151M parameters), but is clearly outperformed by the largest VQAScore. Regarding the two directions, CLIPTROHN-Img is best on I2T but lags behind both VQAScore models in T2I. Given the large difference in size, it is not clear whether the model architecture of VQAScore is key for the high performance.
On the other hand, CLIPTROHN-Text is the best contrastive model for SugarCrepe, outperforming NegCLIP by 8 points and closing the gap with humans to 5 points. However, this strong performance does not reflect on BiVLC, where CLIPTROHN-Text is even worse than CLIPCOCO, NegCLIP and GNM. This highlights our previous finding about the lack of correlation between SugarCrepe and BiVLC results (Section 4).
To put it into the context of the state-of-the-art, the strongest models for SugarCrepe are VQAScore-XXL [Lin et al., 2024] (93.72), CapPa [Tschannen et al., 2024] (92.88) and GPT-4V [Achiam et al., 2023] (92.19) [^2] Results obtained from https://github.com/RAIVNLab/sugar-crepe/tree/main/gpt-4v-results. CLIPTROHN-Text is on par with VQAScore-XXL and outperforms CapPa and GPT-4V, showing it is very strong for image-to-text retrieval, despite its significantly smaller size.
The lower performance of CLIPTROHN-Img on SugarCrepe can be attributed to TROHN-Img containing 10 times less hard negative captions than TROHN-Text, and could improve with more instances.
@@
Which category is the most difficult?
Table 4 shows results of all multimodal models in BiVLC according to caption generation categories. We find that, similar to SugarCrepe, the Swap category is the hardest one for all models also for BiVLC. The group scores for all models are more than 20 points below the scores for the Replace category. But contrary to SugarCrepe, we observe that the Add category is also challenging for models in BiVLC. It is surprising to see the low performance of Open CapPa for that category, since its performance in SugarCrepe is close to perfection. This highlights again the importance of language biases and evaluation protocols that mitigate the effects of those biases. In that sense, it is also interesting to see that CLIPTROHN-Img is the best model for the Add category, outperforming VQAScore-XXL by almost 10 points, despite its significantly smaller size.
Table 4: Results of multimodal models for BiVLC, divided into Replace, Swap and Add categories. Bold for best, underline for second best.
Model Replace Swap Add
I2T T2I Group I2T T2I Group I2T T2I Group
Random 25.00 25.00 16.67 25.00 25.00 16.67 25.00 25.00 16.67
CLIP 82.09 60.36 57.27 46.52 16.16 13.65 70.32 44.63 39.58
CLIPCOCO 87.99 72.42 69.94 51.53 25.07 20.89 83.16 55.58 51.58
NegCLIP 86.37 70.80 68.03 47.08 24.23 18.66 81.26 51.37 48.00
GNM 86.14 68.70 65.75 50.70 17.83 15.88 83.16 58.74 55.37
CLIPTROHN-Text 83.61 70.03 65.84 54.32 25.35 20.61 72.21 55.37 48.42
CLIPTROHN-Img 91.62 79.18 76.61 62.40 32.03 27.86 94.74 69.47 68.00
Open CapPa 70.32 62.79 53.03 40.95 30.64 20.61 14.74 46.32 9.263
VQAScore-XL 85.95 84.18 78.28 65.18 58.77 48.75 77.05 56.63 50.74
VQAScore-XXL 89.85 88.33 83.85 72.14 66.30 56.82 80.42 65.47 58.74
@@
We can also evaluate the detectors on the BiVLC retrieval tasks: (i) for image-to-text retrieval, we first use the visual encoder to predict the type of image (natural or synthetic); if the detection is correct, the output of the text detector is used to see which of the captions has the highest probability to be of the same type as the image, and we select it as the matching caption; (ii) for text-to-image retrieval, we follow the same procedure, but using first the text detector over the given caption, and afterwards the visual detector for the two candidate images. The rightmost columns in Figure 5 show that BiVLC group score is very low for both detectors, underperforming clearly the multimodal models (cf. Table 3). This means that our models are actually learning much more than just distinguishing synthetic and natural images-captions. We already had another evidence of that in Table 3, where all models do equally well on Ipos2T and Ineg2T, i.e. they work equally well for both natural and synthetic images.
@@ 6 Conclusions
We have presented BiVLC, a dataset for bidirectional vision-language compositionality. Our dataset offers the possibility of evaluating the compositional representation of multimodal models in both retrieval directions: image-to-text and text-to-image. We have shown that including text-to-image retrieval leads to novel findings, such as the unbalanced performance of multimodal models in the two retrieval directions, in contrast to humans. We also uncovered the low performance for BiVLC of the models which have been specifically trained for image-to-text retrieval. Furthermore, we have proposed a new way to train models with hard negative images, resulting on the best contrastive model for BiVLC so far. The gap with human performance shows that BiVLC can be used to measure further advances in modelling VLC.
@@ Acknowledgements
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix A.1
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix A.2
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We comply with the ethical codes indicated in the guideline.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See all the links in https://imirandam.github.io/BiVLC_project_page or Appendix E
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix E
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Error bars are not shown because it would be too computationally expensive and time-consuming. You can see computational cost in Appendix E.4
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix E.4
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Appendix E.1
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix E.1
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See the Appendix  B
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix E.1. Since the source data we use has open licenses for non-commercial use, it was not necessary to obtain consent.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix E.1
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix H.1
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Appendix H.2
@@ A.1 Limitations
BiVLC offers captions only in English. It would be interesting to extend the dataset to other languages, as some recent works in vision-language models are already doing [Pouget et al., 2024, Chen et al., 2023, Bugliarello et al., 2022]. Moreover, we only trained contrastive models, due to their suitability for image-to-text and text-to-image retrieval tasks and their availability. In the future, generative multimodal models, which we evaluated but not fine-tuned, could also be explored. Indeed, our approach to fine-tune contrastive models with hard negative images also has its limitations: we evaluated the models in Winoground [Thrush et al., 2022] and we saw that improvements are modest. There are several hypotheses to explain those results and one of them points to the effect of using synthetic images. Deeper analyses are needed to elucidate the real effect of synthetic images to train multimodal models. Finally, as we rely on SugarCrepe  hard negative captions, we also use the same categories. Adding more diversity by extending BiVLC to other categories could be beneficial.
@@ A.2 Societal impacts
Appendix B BiVLC dataset information
Dataset documentation
  • image: COCO 2017 validation image.
  • caption: COCO 2017 validation text describing the COCO image.
  • negative_caption: Negative caption generated from the COCO text description by SugarCrepe [Hsieh et al., 2024].
  • negative_image: Negative image generated from the negative caption by us for BiVLC.
  • type: Category of the negative instances: Replace, Swap or Add.
Figure 5: Example of a BiVLC instance after loading the dataset.
Maintenance plan
Licensing
Author statement
Appendix C TROHN-Text dataset details
@@
Appendix D Detailed evaluation metrics
@@ E.3 Software information
Fine-tuning hyperparameters
@@
We base CLIP fine-tuning on OpenCLIP [Cherti et al., 2022]. Detailed hyperparameters:
@@
We also trained CLIPDet and CLIPTROHN-Img/Det detectors for binary classification by keeping the encoders frozen and adding a sigmoid neuron over the CLS embedding for the image encoder and over the EOT embedding for the text encoder. Detailed hyperparameters:
@@
  • Epochs: We trained the text detector over 10 epochs and the image detectors over 1 epoch. We used validation accuracy as the model selection criterion, i.e. we selected the model with the highest accuracy in the corresponding validation set.
@@
For contrastive models, we base our evaluation on OpenCLIP [Cherti et al., 2022]. We follow all the default hyperparameters used to evaluate models, making sure that when loading the checkpoints we are using QuickGELU as in the base pretrained model [^22] Evaluation bug when using GELU vs QuickGELU https://github.com/RAIVNLab/sugar-crepe/issues/7. For the generative models, we have used the official GitHub repositories of each model and followed the instructions and relied on the evaluation codes provided by their authors.
@@ E.4 Hardware information
Evaluation
All evaluations have been performed on one NVIDIA A100-SXM4-80GB GPU.
@@ Appendix F SugarCrepe templates
Appendix G BiVLC instance examples
@@ Appendix H Crowdsourcing
H.1 Instructions for the different stages
@@
The instructions for this stage are divided into two: image-to-text retrieval (I2T) and text-to-image retrieval (T2I). Both are detailed below, and examples of question as seen by the annotators are depicted in Figures 9 and 10. The obtained human performance for each task can be seen in Table 2. {spverbatim} "Each survey consists of 100 questions/instances. We split the instances into 2 parts (50/50), in which the task changes.
@@
H.2 Hourly wage paid to participants and total amount spent
@@
H.3 Inter-tagger agreement